OpenStudy (anonymous):

Rectangle inside of a rectangle. Inner rectangle: L=2x & W=x There is 7 ft of space between the inner rectangle and the outer rectangle on both sides. Outer Rectangle: L=2x+14 & W=x+14 The area of the outer rectangle is 952ft^2 My job is to find the dimensions of the inner rectangle. Below is the work that I have so far. My mind is not on track for some reason & I keep getting stuck.. please help.

OpenStudy (anonymous):

(2x+14)(x+14) 2x^2+28x+14x+196 2x^2+42x+196 =952 -952 -952 2x^2+42x -756 =0?? (2x+7)(x+7) = 952 2x^2+14x+7x+49 2x^2+21x+49=952 -952 2x^2+21x -903

OpenStudy (anonymous):

your hot

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

like reallllllllly

OpenStudy (radar):

How can the area of both rectangles be 952 sq ft? Isn't one a lot larger than the other?

OpenStudy (anonymous):

The smaller one is inside of the other. That's why I put both, technically its the area of the larger one.

OpenStudy (radar):

From your first equation x =\[-21\pm \sqrt{6069}\over 4\]

OpenStudy (radar):

I believe it is an irrational number but real, using the positive value.